# Object Augmentation Algorithm: Computing virtual object motion and object induced interaction wrench from optical markers

### v2

@@ Abstract
This study addresses the critical need for diverse and comprehensive data focused on human arm joint torques while performing activities of daily living (ADL). Previous studies have often overlooked the influence of objects on joint torques during ADL, resulting in limited datasets for analysis. To address this gap, we propose an Object Augmentation Algorithm (OAA) capable of augmenting existing marker-based databases with virtual object motions and object-induced joint torque estimations. The OAA consists of five phases: (1) computing hand coordinate systems from optical markers, (2) characterising object movements with virtual markers, (3) calculating object motions through inverse kinematics (IK), (4) determining the wrench necessary for prescribed object motion using inverse dynamics (ID), and (5) computing joint torques resulting from object manipulation. The algorithm’s accuracy is validated through trajectory tracking and torque analysis on a 7+4 degree of freedom (DoF) robotic hand-arm system, manipulating three unique objects. The results show that the OAA can accurately and precisely estimate 6 DoF object motion and object-induced joint torques. Correlations between computed and measured quantities were $>$ 0.99 for object trajectories and $>$ 0.93 for joint torques. The OAA was further shown to be robust to variations in the number and placement of input markers, which are expected between databases. Differences between repeated experiments were minor but significant ($p<0.05$). The algorithm expands the scope of available data and facilitates more comprehensive analyses of human-object interaction dynamics.
@@ I Introduction
A marked example of this scarcity is the lack of diverse data on human arm joint torques during activities of daily living (ADL), with the authors being only aware of a single study [6] covering a limited set of ADL. Over the past decades, their data has been consistently used for various fields such as rehabilitative robotics [7, 8], data-driven design optimisation [9, 10], and continual biomechanical investigations of ADL manipulation [11, 12]. The manipulation of objects forms an integral task of ADL, where object manipulation-induced joint torques are at times multiple times higher than limb-related terms. However, the influence of objects was neglected in ADL joint torque estimations of [6]. The lack of representative ADL joint kinetics led the authors of Damerla et al. [13] and Toedtheide et al. [14] to dimension their upper limb prosthetic mechanisms based on the peak human torque capabilities reported in [15]. However, the maximal torques generated by human joints exceed functional requirements, resulting in excessively heavy prosthetic devices — a significant concern for users. [16, 17, 18].
@@
Ulna/Radial deviation
@@ II-A1 Grasping experiments
A 4 DoF transhumeral prosthesis platform, the Artificial Neuromuscular Prosthesis (ANP)[14] equipped with a 7 DoF robotic hand (IH2 Azzurra, Prensilia SRL, Italy) was used to perform grasping experiments. The kinematics of the ANP and IH2 reflect those of the human elbow, forearm, wrist and hand. The objects utilized were a 425 ml tin can, a 500 ml water bottle, and a cordless drill weighing 720 g, Bosch GSR 12V-35 without battery. They were manually placed in the open robotic hand and grasped with the pre-programmed cylindrical grasp. Fig 1. depicts corresponding joint configuration trajectories, where the ANP was mounted on a static fixture at the elbow joint. Subsequently, objects were moved by the ANP along four pre-programmed motions as follows:
@@ II-B Object Augmentation Algorithm
    virtual object kinematics
@@
  • Optical markers: trajectories of the individual markers
  • Object: dimensions, mass, and inertia $\in\{D,m,I\}$.
@@
$\{M_{w},M_{h}\}$ optical wrist and hand marker trajectories
@@
$\alpha$ wrist deviation angle
@@
Phase (1) computed the robotic hand coordinate systems $F_{hand,i}\in\{\vec{n_{i}},\vec{p_{i}},\vec{f_{i}}\}$ for each frame $i$ of input markers $\{M_{w},M_{h}\}$, as shown in Fig. 3. The first vector $\vec{n}$ represents the normal vector of the palm. It was calculated using the Singular Value Decomposition (SVD) method applied to $M_{h}$, where $\vec{n}$ was found as the right singular vector associated with the smallest singular value, equation (2). The intended direction of $\vec{n}$ depicted in Fig. 2 is not guaranteed by SVD. The direction of $\vec{n_{i}}$ reverses if there is an obtuse angle between $\vec{n_{i}}$ and the result of the cross product between the vector $\vec{M_{w}}$ and $M_{w,k}-M_{h,k}$ (the vector from any wrist to any hand marker, $k$ = arbitrary index) as shown in equation (2)
@@
In phase (2), object movements were characterized by computing trajectories of virtual markers placed on the virtual objects $V_{obj}\in\{V_{can},V_{bottle},V_{drill}\}$. $V_{can}$ and $V_{bottle}$ were modelled as a cylinder with a radius $r$, height $h$, grasped by the robotic hand with a palm thickness of $H$, while the drill was modelled as cylinder resting on a cube, see Fig. 1. Three virtual markers (VM) were placed on each of the $V_{obj}$, one on top centre of the cylinder ($VM_{t}$), the second on the Center of Mass (CoM) of ($VM_{c}$), and the third on the cylinder circumference at the height of the CoM ($VM_{s}$) of the $V_{obj}$, as shown in Fig. 1. The placement of the aforementioned virtual markers was according to $F_{hand}$ frame as described by (4), with VM coordinates of $VM_{t}$ and $VM_{c}$ collected in table I.
@@
Virtual objects, $V_{obj}$
Phase (3), calculates object motions $P_{obj}$ through Inverse Kinematics (IK) (OpenSim 4.4 IK solver [27]), where virtual marker trajectories $M_{o}$ were subjected to a 3rd order low pass filter (cutoff 6 Hz). Resulting motions were either $\in\mathbb{R}^{6}$ or span a subspace if environmental constraints exist (e.g., partial support by surfaces, joints, etc.). Phase (4), calculates the wrench $W_{obj}\in\mathbb{R}^{6}$ necessary to move objects along the prescribed motion by ID (OpenSim 4.4 ID tool) as well as dynamic object models. Object models were based on the real objects described in subsection II-A1. The mass and inertia properties of the $V_{obj}$ (table I) are reflective of the physical objects (can, bottle, and drill).
@@ II-C Validation of OAA and statistical analysis
The validation of the OAA includes a comparison of measured object trajectories and joint torques from ANP with that of estimated by OAA for the following scenarios:
@@
The average hand marker centroid distance for each $\lvert M_{h}\rvert$ was produced by changing the location of each marker in the $M_{h}$ set a 1000 times per marker, hence 5000 centroids were generated. The marker placement was randomised within a diameter of 6 cm, corresponding to the average breadth of the female hand [28] minus one marker diameter (1.2 cm) while ensuring a minimal inter-marker distance of 3 cm.
@@ III-B Object trajectory tracking
Table II collects error statistics for all markers ($VM$, $VM_{off}$ vs ground truth) averaged over three repetitions per object. Fig. 5 (a) shows marker trajectory errors as L2 distances between marker coordinates between virtual and ground truth.
@@ III-C Object induced joint torques
The maximum marker errors reported by OpenSim for IK were 92$\pm$36 $\mu$m for virtual and 118$\pm$19 $\mu$m for ground truth markers, averaged over all trials. Joint torques reported below include prosthesis and object manipulation wrench-induced terms according to equation (1). Joints are labelled EF for the elbow flexion/extension, PS for forearm pronation/supination, WF for wrist flexion/extension, and WD for wrist Ulnar/Radial deviation. Torques were compared for cases:
@@ 3. III
Fig. 6 reports joint torque distributions for cases I-IV. Black boxes indicate case I (measured) and coloured boxes case II (computed) joint torques, whereas brown boxes relate to simulated prosthesis torques in the absence of objects. Grey-scale boxes describe torque terms produced solely by the objects described in cases II-IV. Case I and II torques were similarly distributed for all joints except PS. In PS, percentile and peak torques were higher in the computed case. In EF and WD joints, measured peak torques were higher than computed torques, while they were equal in the WF joint. Distributions of $\boldsymbol{\tau}_{Vobj}$, $\boldsymbol{\tau}_{Vobj\_off}$, and $\boldsymbol{\tau}_{Gobj\_off}$ were indistinguishable for all joints and objects. The ratio between $\boldsymbol{\tau}_{P\_s}$ and object-induced torques showed different qualities between joints. In EF, $\boldsymbol{\tau}_{P\_s}$ was higher than object terms, while they were lower in PS. For wrist joints, $\boldsymbol{\tau}_{P\_s}$ and object terms were similarly distributed.
@@
Figs. 7 (b) and (c) compare Case I (grey-scale, dashed) and Case II (coloured) torque trajectories during the first repetition. Additional dotted lines mark isolated prosthesis-related torque terms for the same cases. Case I $\boldsymbol{\tau}_{P}$ was obtained in physical experiments without objects, and Case II $\boldsymbol{\tau}_{P_{s}}$ in simulation by (1) without object wrench. Computed and measured torque trajectories correlated well, except for wrist WF and WD joints during their respective motions. There $\boldsymbol{\tau}_{P_{s}}$cong 0 Nm diverged from measured $\boldsymbol{\tau}_{P}$, with the latter exhibiting similar trajectories as in trials including objects. Fast elbow motion produced torque spikes in the EF and WD joints.
TABLE III: Joint Torque statistics - measured against computed, and percentile torque contributions of objects averaged over three repetitions per trial. Errors are in Nm.
@@
Table III reports statistical values of Case I against Case II errors, averaged over three repetitions per trial. Standard deviations are omitted due to their small magnitudes. In comparison to peak torques plotted in Figs 7 (b) and (c), average and RMS torque errors are low, while maximal errors reach 10-20% of peak magnitudes. Pearson correlation is high at 0.93 to 0.98 for all joints and objects ($p<<0$). Percentile object torques were between 34% and 80% of limb torques, where influences were generally higher for 50th percentile torques. Percentiles were computed as limb percentile over object percentile torques.
@@ IV Discussion
The performance of the OAA was experimentally validated on the 4DoF ANP platform [14], equipped with accurate torque sensors in each joint and using a gold standard optical motion capture system. Three unique objects were grasped with a 7 DoF robotic anthropomorphic hand. The system was moved along 4 pre-programmed joint motions. Motions were designed to test each joint in isolation. The experimental protocol involved three repetitions per object, with objects being re-grasped between each repetition. Grasping manoeuvres were executed without strict adherence to maintaining consistent relative configurations between the object and the hand. Object selection was deliberate, aiming to expose the limitations of the algorithm: the diameter of the Can was at the limits of the robotic hand-grasping capabilities; the elongated shape of the bottle accentuated tracking errors of $VM_{t}$; and the forward-facing geometry of the drill led to potential misalignment issues around the $\vec{p}$ axis.
@@
Virtual markers were positioned within hand coordinate frames through transformations relying exclusively on object dimensions. This approach mirrors real-world scenarios where object trajectories are unavailable and grasping procedures lack sufficient documentation for precise replication. Virtual object trajectories exhibited a high correlation with motion-captured ground truth trajectories, validating the accuracy of the proposed method. OpenSim guidelines recommend maximal and root mean square (RMS) marker tracking errors to be below 20 mm and 10 mm, respectively, for static poses. Our findings consistently fell well below these thresholds during static and dynamic motions. One exception occurred with the Drill’s virtual marker trajectories ($VM_{t}$) during supination, attributed to the neglected rotation around the $\vec{p}$ axis for the Drill. Wrist deviation phases were performed with the forearm pronated at 90, alleviating torque limitations of the ANPs wrist by reducing gravitational joint loads. Forearm pronation generally increased marker errors and was the reason for increased tracking errors during wrist deviation phases. This led to relatively large error deviations for both can and drill $VM_{t}$s, reflecting grip variations expected in trials with human subjects.
@@
We successfully employed the OAA to augment motions pertaining to object manipulation of the ADL Human Arm Motion Dataset [32]. Joint quantities were computed for a wide range of objects, enabled by the OAA’s ability to compute prescribed motions and induced joint torques for arbitrary objects. In addition, the OAA allowed the simulation of joint torques produced by small or kinematically constrained items, such as turning a key in a lock and opening a door. Respective results are reported in [33]. Other databases, such as [23, 24] may be augmented similarly, enriching the information on ADL joint kinetics available in literature.
@@ V Appendix
An open source implementation of the OAA is available at https://github.com/ChristopherHerneth/ObjectAugmentationAlgorithm/tree/main. Accompanying video material may be found here https://youtu.be/8oz-awvyNRA. This document is the Author’s Original Manuscript. For the final version please refer to the proceedings of IROS 2024.

### v3

@@ Abstract
This study addresses the critical need for diverse and comprehensive data focused on human arm joint torques while performing activities of daily living (ADL). Previous studies have often overlooked the influence of objects on joint torques during ADL, resulting in limited datasets for analysis. To address this gap, we propose an Object Augmentation Algorithm (OAA) capable of augmenting existing marker-based databases with virtual object motions and object-induced joint torque estimations. The OAA consists of five phases: (1) computing hand coordinate systems from optical markers, (2) characterising object movements with virtual markers, (3) calculating object motions through inverse kinematics (IK), (4) determining the wrench necessary for prescribed object motion using inverse dynamics (ID), and (5) computing joint torques resulting from object manipulation. The algorithm’s accuracy is validated through trajectory tracking and torque analysis on a 5+4 degree of freedom (DoF) robotic hand-arm system, manipulating three unique objects. The results show that the OAA can accurately and precisely estimate 6 DoF object motion and object-induced joint torques. Correlations between computed and measured quantities were $>$ 0.99 for object trajectories and $>$ 0.93 for joint torques. The OAA was further shown to be robust to variations in the number and placement of input markers, which are expected between databases. Differences between repeated experiments were minor but significant ($p<0.05$). The algorithm expands the scope of available data and facilitates more comprehensive analyses of human-object interaction dynamics.
@@ I Introduction
A marked example of this scarcity is the lack of diverse data on human arm joint torques during activities of daily living (ADL), with the authors being only aware of a single study [6] covering a limited set of ADL. Over the past decades, their data has been consistently used for various fields such as rehabilitative robotics [7, 8], data-driven design optimisation [9, 10], and continual biomechanical investigations of ADL manipulation [11, 12]. The manipulation of objects forms an integral task of ADL, where object manipulation-induced joint torques are at times multiple times higher than limb-related terms. However, the influence of objects was neglected in ADL joint torque estimations of [6]. The lack of representative ADL joint kinetics led the authors of Damerla et al. [13] and Toedtheide et al. [14] to dimension their upper limb prosthetic mechanisms based on the peak human torque capabilities reported in [15]. However, the maximal torques generated by human joints exceed functional requirements, resulting in excessively heavy prosthetic devices — a significant concern for users [16, 17, 18].
@@
Ulnar/Radial deviation
@@ II-A1 Grasping experiments
A 4 DoF transhumeral prosthesis platform, the Artificial Neuromuscular Prosthesis (ANP)[14] equipped with a 5 DoF robotic hand (IH2 Azzurra, Prensilia SRL, Italy) was used to perform grasping experiments. The kinematics of the ANP and IH2 reflect those of the human elbow, forearm, wrist and hand. The objects utilized were a 425 ml tin can, a 500 ml water bottle, and a cordless drill weighing 720 g, Bosch GSR 12V-35 without battery. They were manually placed in the open robotic hand and grasped with the pre-programmed cylindrical grasp. Fig 1. depicts corresponding joint configuration trajectories, where the ANP was mounted on a static fixture at the elbow joint. Subsequently, objects were moved by the ANP along four pre-programmed motions as follows:
@@ II-B Object Augmentation Algorithm
    virtual object kinematics by Inverse Kinematics (IK)
@@
  • Optical markers: trajectories of markers $\in$ $M_{w}$, and $M_{h}$.
  • Object: dimensions, mass, and inertia $\in\{D,m,I\}$, and a set of virtual object markers and their location within the hand.
@@
$\{M_{w},M_{h}\}$ optical wrist and hand marker frames
@@
$\alpha$ wrist deviation angle trajectory
@@
Phase (1) computed the robotic hand coordinate systems $F_{hand,i}\in\{\vec{n_{i}},\vec{p_{i}},\vec{f_{i}}\}$ for each frame $i$ of input markers $\{M_{w},M_{h}\}$, as shown in Fig. 3. The first vector $\vec{n}$ represents the normal vector of the palm. It was calculated using the Singular Value Decomposition (SVD) method applied to $M_{h}$, where $\vec{n}$ was found as the right singular vector associated with the smallest singular value, equation (2). The intended direction of $\vec{n}$ depicted in Fig. 2 is not guaranteed by SVD. The direction of $\vec{n_{i}}$ reverses if there is an obtuse angle between $\vec{n_{i}}$ and the result of the cross product between the vector $\vec{M_{w,i}}$ (unit vector from the Ulnar to the Radial styloid) and $M_{h,i,k}-M_{w,i,k}$ (the vector from any wrist to any hand marker, $k$ = arbitrary index) as shown in equation (2)
@@
In phase (2), object movements were characterized by computing trajectories of virtual markers placed on the virtual objects $V_{obj}\in\{V_{can},V_{bottle},V_{drill}\}$. $V_{can}$ and $V_{bottle}$ were modelled as a cylinder with a radius $r$, height $h$, grasped by the robotic hand with a palm thickness of $H$, while the drill was modelled as cylinder resting on a cube, see Fig. 1. Three virtual markers (VM) were placed on each of the $V_{obj}$, one on top centre of the cylinder ($VM_{t}$), the second on the object’s Center of Mass (CoM) ($VM_{c}$), and the third on the cylinder’s circumference at the height of the CoM ($VM_{s}$), as shown in Fig. 1. The placement of the aforementioned virtual markers was according to $F_{hand}$ frame as described by (4), with VM coordinates of $VM_{t}$ and $VM_{c}$ collected in table I.
@@
Phase (3), calculated object motions $P_{obj}$ through IK (OpenSim 4.4 IK solver [27]), where virtual marker trajectories $M_{o}$ were subjected to a 3rd order low pass filter (cutoff 6 Hz). Resulting motions were either $\in\mathbb{R}^{6}$ or span a subspace if environmental constraints exist (e.g., partial support by surfaces, joints, etc.).
Phase (4), calculated the wrench $W_{obj}\in\mathbb{R}^{6}$ necessary to move objects along the prescribed motion by ID (OpenSim 4.4 ID tool) as well as dynamic object models. Object models were based on the real objects described in subsection II-A1, with mass and inertia properties collected in table I.
@@ II-C Validation of OAA and statistical analysis
The validation of the OAA includes a comparison of measured object trajectories and joint torques from the ANP with that of estimated by the OAA for the following scenarios:
@@
The average hand marker centroid distance for each $\lvert M_{h}\rvert$ was produced by changing the location of markers in the $M_{h}$ set a 1000 times, hence 4000 centroids were generated. The marker placement was randomised within a diameter of 6 cm, corresponding to the average breadth of the female hand [28] minus one marker diameter (1.2 cm) while ensuring a minimal inter-marker distance of 3 cm.
@@ III-B Object trajectory tracking
Table II collects error statistics markers ($VM$, $VM_{off}$ vs ground truth) averaged over three repetitions per object. Fig. 5 (a) shows marker trajectory errors as L2 distances between virtual and ground truth marker coordinates.
@@ III-C Object induced joint torques
The maximum marker errors reported by OpenSim for IK were 92$\pm$36 $\mu$m for virtual and 118$\pm$19 $\mu$m for ground truth markers, averaged over all trials. Joint torques reported below include prosthesis and object manipulation wrench-induced terms according to equation (1). Joints are labeled EF for the elbow flexion/extension, PS for forearm pronation/supination, WF for wrist flexion/extension, and WD for wrist Ulnar/Radial deviation. Torques were compared for cases:
@@ 3. III
Fig. 6 reports joint torque distributions for cases I-IV. Black boxes indicate case I (measured) and coloured boxes case II (computed) joint torques, whereas brown boxes relate to simulated prosthesis torques in the absence of objects. Grey-scale boxes describe torque terms produced solely by the objects described in cases II-IV. Case I and II torques were similarly distributed for all joints except PS. In PS, the percentile and peak torques were higher in the computed case. In EF and WD joints, measured peak torques were higher than computed torques, while they were equal in the WF joint. Distributions of $\boldsymbol{\tau}_{Vobj}$, $\boldsymbol{\tau}_{Vobj\_off}$, and $\boldsymbol{\tau}_{Gobj\_off}$ were indistinguishable for all joints and objects. The ratio between $\boldsymbol{\tau}_{P\_s}$ and object-induced torques showed different qualities between joints. In EF, $\boldsymbol{\tau}_{P\_s}$ was higher than object terms, while they were lower in PS. For wrist joints, $\boldsymbol{\tau}_{P\_s}$ and object terms were distributed similarly.
@@
Figs. 7 (b) and (c) compare Case I (grey-scale, dashed) and Case II (coloured) torque trajectories during the first repetition. Additional dotted lines mark prosthesis-related torque terms for the same cases. Case I $\boldsymbol{\tau}_{P}$ was obtained in physical experiments without objects, and Case II $\boldsymbol{\tau}_{P_{s}}$ in simulation by (1) without object wrench. Computed and measured torque trajectories correlated well, except for wrist WF and WD joints during their respective motions. There $\boldsymbol{\tau}_{P_{s}}$cong 0 Nm diverged from measured $\boldsymbol{\tau}_{P}$, with the latter exhibiting similar trajectories as in trials including objects. Fast elbow motion produced torque spikes in the EF and WD joints.
TABLE III: Joint Torque statistics - measured against computed, and percentile torque contributions of objects averaged over three repetitions per trial. Torques are in Nm.
@@
Table III reports statistical values of Case I against Case II errors, averaged over three repetitions per trial. Standard deviations are omitted due to their small magnitudes. In comparison to peak torques plotted in Figs 7 (b) and (c), average and RMS torque errors are low, while maximal errors reach 10-20% of peak magnitudes. Pearson correlation is high at 0.93 to 0.98 for all joints and objects ($p<<0$). Percentile object torques were between 34% and 80% of limb torques, where influences were generally higher for 50th percentile torques. Percentages were computed as limb percentile over object percentile torques.
@@ IV Discussion
The performance of the OAA was experimentally validated on the 4DoF ANP platform [14], equipped with accurate torque sensors in each joint and using a gold standard optical motion capture system. Three unique objects were grasped with a 5 DoF robotic anthropomorphic hand. The system was moved along 4 pre-programmed joint motions. Motions were designed to test each joint in isolation. The experimental protocol involved three repetitions per object, with objects being re-grasped between each repetition. Grasping manoeuvres were executed without strict adherence to maintaining consistent relative configurations between the object and the hand. Object selection was deliberate, aiming to expose the limitations of the algorithm: the diameter of the Can was at the limits of the robotic hand-grasping capabilities; the elongated shape of the bottle accentuated tracking errors of $VM_{t}$; and the forward-facing geometry of the drill led to potential misalignment issues around the $\vec{p}$ axis.
@@
Virtual markers were positioned within hand coordinate frames through transformations relying exclusively on object dimensions. This approach mirrors real-world scenarios where object trajectories are unavailable and grasping procedures lack sufficient documentation for precise replication. OpenSim guidelines recommend maximal and root mean square (RMS) marker tracking errors to be below 20 mm and 10 mm, respectively, for static poses. Our findings consistently fell well below these thresholds during static and dynamic motions. One exception occurred with the Drill’s virtual marker trajectories ($VM_{t}$) during supination, attributed to the neglected rotation around the $\vec{p}$ axis for the Drill. Virtual object trajectories exhibited a high correlation with motion-captured ground truth trajectories. Wrist deviation phases were performed with the forearm pronated at 90, alleviating torque limitations of the ANPs wrist by reducing gravitational joint loads. Forearm pronation generally increased marker errors and was the reason for increased tracking errors during wrist deviation phases. This led to relatively large error deviations for both can and drill $VM_{t}$s, reflecting grip variations expected in trials with human subjects.
@@
We successfully employed the OAA to augment motions pertaining to object manipulation of the ADL Human Arm Motion Dataset [32]. Joint quantities were computed for a wide range of objects, enabled by the OAA’s ability to compute prescribed motions and induced joint torques for arbitrary objects. In addition, the OAA allowed the simulation of joint torques produced by small or kinematically constrained items, such as turning a key in a lock and opening a door. Respective results are presented in [33]. Other databases, such as [23, 24] may be augmented similarly, enriching the information on ADL joint kinetics available in literature.
@@ V Appendix
An open source implementation of the OAA is available at https://github.com/ChristopherHerneth/ObjectAugmentationAlgorithm/tree/main. Accompanying video material may be found here https://youtu.be/8oz-awvyNRA.